# PDEformer: Towards a Foundation Model for One-Dimensional Partial Differential Equations

### v1

@@ Abstract
This paper introduces PDEformer, a neural solver for partial differential equations (PDEs) capable of simultaneously addressing various types of PDEs. We advocate representing the PDE in the form of a computational graph, facilitating the seamless integration of both symbolic and numerical information inherent in a PDE. A graph Transformer and an implicit neural representation (INR) are employed to generate mesh-free predicted solutions. Following pretraining on data exhibiting a certain level of diversity, our model achieves zero-shot accuracies on benchmark datasets that surpass those of adequately trained expert models. Additionally, PDEformer demonstrates promising results in the inverse problem of PDE coefficient recovery.
@@ 1 Introduction and Related Work
Drawing inspirations from successful experiences in natural language processing and computer vision, we aim to develop a foundation PDE model with the highest generality, capable of handling any PDE in the ideal case. Given a new PDE to be solved, we only need to make a direct (zero-shot) inference using this model, or fine-tune it only for a few steps using a relatively small number of solution snapshots. By leveraging the power of generality, its solution accuracy is anticipated to surpass expert models that are adequately trained to solve a specific type of PDE. Besides, such a foundation model should be easily adapted to tackle with down-stream tasks, including inverse problems, inverse design, optimal control, etc.
A PDE to be solved would involve two parts of information: one is the symbolic part specifying the mathematical form of the PDE, and the other is the numeric part that includes the PDE coefficients, initial and boundary values, etc. Typical neural operators like Fourier Neural Operator (FNO) (Li et al., 2021) and DeepONet (Lu et al., 2021) deal with a specific form of PDE, and only need to take the numeric information as the network input. However, in order to construct a foundation model generalizable to different PDEs, the symbolic information has to be integrated seamlessly.
@@ 3.2 Forward problem
The pretrained PDEformer is highly versatile in handling various equations. Its performance on forward problems is evaluated using parametric PDEs from PDEBench (Takamoto et al., 2022), including Burgers’, Advection, and 1D Reaction-Diffusion PDEs. Comparative analysis is conducted with neural operator models tailored for individual PDEs. In this context, neural operators receive initial conditions and predict the entire solution field. Notably, all previous methods are trained and tested separately on different datasets, while the PDEformer model shows zero-shot inference across all test datasets post pretraining. Furthermore, the PDEformer-FT model involves an additional fine-tuning process on the corresponding PDEBench dataset.
@@
In Table 1, the pretrained PDEformer model showcases zero-shot proficiency by attaining reasonable accuracy in all in-distribution tests [^2] Here, “in-distribution” and “out-of-distribution”(OoD) refer to the range of the PDE coefficients. In-distribution PDEs have coefficients lying within the range of the pretraining data, whereas OoD samples fall outside this range. Note that PDEBench datasets labeled as “in-distribution” are not utilized during pretraining. Therefore, we term the corresponding PDEformer inference as “zero-shot”. . Remarkably, for the Burgers’ equation with $\nu=0.1$ and $0.01$, the zero-shot PDEformer outperforms the sufficiently trained FNO model. This superior performance stems from the pretrained PDEformer’s exposure to diverse PDEs during pretraining, from which the model learns a generalizable law to outperform models trained specifically for individual PDEs. The fine-tuned PDEformer consistently excels in all tests, further highlighting the robustness and versatility of our approach in solving a wide range of PDEs.
@@
Table 1: Test relative $L^{2}$ error on PDEBench. Format the first and second best outcomes in bold and underline respectively.
Model Burgers Adection Reaction-Diffusion
$\nu=0.1$ $\nu=0.01$ $\nu=0.001$ $\beta=0.1$ $\beta=1$ $\nu=1,~{}\rho=1$
U-Net (Ronneberger et al., 2015) 0.1627 0.2253 0.2431 0.0873 0.2655 0.0126
DeepONet (Lu et al., 2021) 0.0699 0.1791 0.2010 0.0186 0.0187 0.0015
FNO (Li et al., 2021) 0.0155 0.0445 0.0700 0.0089 0.0097 0.0018
PDEformer (Ours) 0.0103 0.0309 0.0921 (OoD) 0.0119 0.4000 (OoD) 0.7399 (OoD)
PDEformer-FT (Ours) 0.0046 0.0146 0.0295 0.0043 0.0075 0.0009
@@
To compare the efficiency of fine-tuning PDEformer with training traditional neural operators, we conduct a comparative analysis on the Advection equation ($\beta=1$, OoD) dataset with a limited number of $100$ training samples. As depicted in Figure 3, PDEformer rapidly reaches convergence in about just 100 iterations. Conversely, the FNO model, trained from scratch, takes several thousands of iterations to converge and results in a higher test error. This contrast highlights PDEformer’s swift and accurate adaptability, marking a significant advancement over neural operators.
@@
Figure 3: Comparing the speed of fine-tuning PDEformer with training FNO from scratch. The right subfigure uses a logarithmic scale for the $x$-axis, whereas the left employs a linear scale.
@@ B.2 PDEBench Datasets
In this subsection, we present an overview of three 1D PDE datasets derived from PDEBench, which we employed in our experimental analysis. Each dataset, tailored to a specific PDE type and coefficient configuration, encompasses 10k instances. For our training purposes, we utilized 9k samples from each dataset, reserving the remaining 1k samples for testing. It is crucial to note that every dataset in PDEBench adheres to periodic boundary conditions.
@@
To ensure equitable comparisons among the baseline models, we standardize the resolution of all PDEBench samples to $256\times 256$. More specifically, the original PDEBench datasets have a spatial resolution of $1024$, which is downsampled to $256$. The original number of recorded time-steps is $201$ for the Burgers and Advection datasets and $101$ for the one-dimensional Reaction-Diffusion dataset, and a linear interpolation is utilized to obtain a temporal resolution of $256$. It is important to note that PDEformer makes mesh-free predictions, enabling us to set the temporal resolution to $101$ for the pretraining dataset, and $256$ for the PDEBench dataset. For FNO and U-Net, the initial value is repeated $256$ times to form the two-dimensional data with resolution $256\times 256$, and then taken as the network input.
@@ Appendix D Metric and Detailed Results
In our investigation of the forward problem, we embarked on a detailed investigation to assess the model’s learning efficiency with limited data. Specifically, we reduced the training dataset size from 9k to 100 and 1k samples. As depicted in Figure 8, the fine-tuned PDEformer model notably excels, outperforming all other methods. Moreover, the zero-shot PDEformer establishes a commendably high benchmark, demonstrating robust performance without any fine-tuning. It is particularly noteworthy that under OoD conditions, such as in the Advection $(\beta=1)$ and Reaction-Diffusion scenarios, the fine-tuned PDEformer rapidly attains superior results. This highlights the model’s few-shot learning [^6] In this work, we use the term few-shot learning to describe the model’s proficiency in adapting to and learning from new data that falls outside the distribution of the training set, using only a small number of examples for fine-tuning. ability in adapting to unfamiliar scenarios.

### v3

@@ Abstract
This paper introduces PDEformer, a neural solver for partial differential equations (PDEs) capable of simultaneously addressing various types of PDEs. We propose to represent the PDE in the form of a computational graph, facilitating the seamless integration of both symbolic and numerical information inherent in a PDE. A graph Transformer and an implicit neural representation (INR) are employed to generate mesh-free predicted solutions. Following pretraining on data exhibiting a certain level of diversity, our model achieves zero-shot accuracies on benchmark datasets that is comparable to those of specifically trained expert models. Additionally, PDEformer demonstrates promising results in the inverse problem of PDE coefficient recovery.
@@ 1 Introduction and Related Work
The efficient solution of PDEs plays a crucial role in various scientific and engineering domains, from simulating physical phenomena to optimizing complex systems. In recent years, many learning-based PDE solvers have emerged. Some methods (Raissi et al., 2019; Sirignano & Spiliopoulos, 2018; Ee & Yu, 2017; Zang et al., 2020) represent the approximate PDE solution with a neural network, and are tailored to individual PDEs. Other approaches, such as neural operators like Fourier Neural Operator (FNO) (Li et al., 2021) and DeepONet (Lu et al., 2021), tackle parametric PDEs by taking the PDE parameters (coefficient fields, initial conditions, etc.) as network inputs. While these methods exhibit a higher level of generality, their capability is still limited to solving a specific type of PDE.
Drawing inspirations from successful experiences in natural language processing and computer vision, we aim to develop a foundation PDE model with the highest generality, capable of handling any PDE in the ideal case. Given a new PDE to be solved, we only need to make a direct (zero-shot) inference using this model, or fine-tune it only for a few steps using a relatively small number of solution snapshots. By leveraging the power of generality, foundation models have demonstrated great potential in capturing the similarity inherent in a wide range of tasks, and producing high-quality feature representations that are beneficial to various applications (Bommasani et al., 2022; Zhou et al., 2023). Specific to the realm of scientific computing, we anticipate such a foundation PDE model can achieve high solution accuracy, which is comparable with or even surpass expert models that are trained to solve a specific type of PDE. Besides, it should be easily adapted to tackle with down-stream tasks, including inverse problems, inverse design, optimal control, etc.
A PDE to be solved would involve two parts of information: one is the symbolic part specifying the mathematical form of the PDE, and the other is the numeric part that includes the PDE coefficients, initial and boundary values, etc. Typical neural operators like FNO and DeepONet deal with a specific form of PDE, and only need to take the numeric information as the network input. However, in order to construct a foundation model generalizable to different PDEs, the symbolic information has to be integrated seamlessly.
@@ 3.2 Forward problem
The pretrained PDEformer is highly versatile in handling various equations. Its performance on forward problems is evaluated using parametric PDEs from PDEBench (Takamoto et al., 2022), including Burgers’, Advection, and 1D Reaction-Diffusion PDEs. Comparative analysis is conducted with neural operator models tailored for individual PDEs. In this context, neural operators receive initial conditions and predict the entire solution field. Notably, all previous methods as well as PDEformer-FS are trained from-scratch and tested separately on different datasets, while the PDEformer model shows zero-shot inference across all test datasets post pretraining. Furthermore, the PDEformer-FT model involves an additional fine-tuning process on the corresponding PDEBench dataset.
@@
In Table 1, the pretrained PDEformer model showcases zero-shot proficiency by attaining reasonable accuracy in all in-distribution tests [^2] Here, “in-distribution” and “out-of-distribution”(OoD) refer to the range of the PDE coefficients. In-distribution PDEs have coefficients lying within the range of the pretraining data, whereas OoD samples fall outside this range. Note that PDEBench datasets labeled as “in-distribution” are not utilized during pretraining. Therefore, we term the corresponding PDEformer inference as “zero-shot”. . Remarkably, for Burgers’ equation with $\nu=0.1$ and $0.01$, the zero-shot PDEformer outperforms all the baseline models trained specifically on these datasets. Such superior performance can be partially attributed to the network architecture we have utilized, as PDEformer already exhibits competitive performance when trained from-scratch. We believe that the additional improvement of the pretrained PDEformer stems from its exposure to diverse PDEs during pretraining, from which the model may learn a generalizable law to outperform models trained specifically for individual PDEs. Results of the out-of-distribution tests can be found in Table 3 in the Appendix. The fine-tuned PDEformer consistently excels in all in-distribution and out-of-distribution tests, further highlighting the robustness and versatility of our approach in solving a wide range of PDEs.
@@
Table 1: Test relative $L^{2}$ error on PDEBench, in which the PDE coefficients lie within the range of the pretraining data. We format the first and second best outcomes in bold and underline, respectively.
Model Burgers Advection
$\nu=0.1$ $\nu=0.01$ $\beta=0.1$
U-Net (Ronneberger et al., 2015) 0.1627 0.2253 0.0873
Autoregressive U-Net 0.2164 0.2688 0.0631
DeepONet (Lu et al., 2021) 0.0699 0.1791 0.0186
FNO (Li et al., 2021) 0.0155 0.0445 0.0089
PDEformer-FS (Ours) 0.0135 0.0399 0.0124
PDEformer (Ours) 0.0103 0.0309 0.0119
PDEformer-FT (Ours) 0.0046 0.0146 0.0043
@@
Thanks to the high-quality initialization obtained after pretraining, general foundation models are known to exhibit efficient adaptation to new tasks (Bommasani et al., 2022; Zhou et al., 2023). To compare the efficiency of fine-tuning PDEformer with training traditional expert models, we conduct a comparative analysis on the Advection equation ($\beta=1$, OoD) dataset with a limited number of $100$ training samples. As depicted in Figure 3, PDEformer rapidly reaches convergence in about just 100 iterations. Conversely, the FNO model, trained from scratch, results in a higher test error even after several thousands of iterations. Indeed, it is possible for traditional neural operators to start from a better initialization. However, designed for a specific type of PDE, they cannot be pretrained on 500k data samples containing diverse PDEs as PDEformer does. The valid option left for us is to pretrain them on one different PDE, and then transfer to the target setting, which could be much less efficient. Post pretraining on 9k samples of the Advection equation with $\beta=0.1$ for 1k iterations, the FNO-FT model only exhibits a limited improvement over the corresponding from-scratch version, as can be seen in the figure. This contrast highlights the pretrained PDEformer’s swift and accurate adaptability, marking a significant advancement over existing expert models.
@@
Figure 3: Comparing the speed of fine-tuning PDEformer with training FNO from scratch and fine-tuning a pretrained FNO. The right subfigure uses a logarithmic scale for the $x$-axis, whereas the left employs a linear scale. The vertical lines correspond to $100$ iterations.
@@ 3.3 Inverse problem
Acknowledgments
This work is supported in part by the National Science and Technology Major Project (2022ZD0117804). Bin Dong is supported in part by the New Cornerstone Investigator Program.
@@ B.2 PDEBench Datasets
In this subsection, we present an overview of three 1D PDE datasets derived from PDEBench, which we employed in our experimental analysis. Each dataset, tailored to a specific PDE type and coefficient configuration, encompasses 10k instances. For our training purposes, we utilized 9k samples from each dataset, reserving the remaining 1k samples for testing. It is crucial to note that all these PDEBench datasets adhere to periodic boundary conditions.
@@
To ensure equitable comparisons among the baseline models, we standardize the resolution of all PDEBench samples to $256\times 256$. More specifically, the original PDEBench datasets have a spatial resolution of $1024$, which is downsampled to $256$. The original number of recorded time-steps is $201$ for the Burgers and Advection datasets and $101$ for the one-dimensional Reaction-Diffusion dataset, and a linear interpolation is utilized to obtain a temporal resolution of $256$. It is important to note that PDEformer makes mesh-free predictions, enabling us to set the temporal resolution to $101$ for the pretraining dataset, and $256$ for the PDEBench dataset. For FNO and U-Net (non-autoregressive case), the initial value is repeated $256$ times to form the two-dimensional data with resolution $256\times 256$, and then taken as the network input.
@@ C.3 Training Setting
Autoregressive U-Net
   learning_rate $0.0001$ The initial learning rate for the optimizer
@@ Appendix D Metric and Detailed Results
In our investigation of the forward problem, Table 3 compares the prediction results on PDEs with coefficients lying outside the range of the pretraining data. Note that all baseline methods as well as PDEformer-FS do not involve a pretraining process. We also embarked on a detailed investigation to assess the model’s learning efficiency with limited data. Specifically, we reduced the training dataset size from 9k to 100 and 1k samples. As depicted in Figure 8, the fine-tuned PDEformer model notably excels, outperforming all other methods in the test. Moreover, the zero-shot PDEformer establishes a commendably high benchmark, demonstrating robust performance without any fine-tuning. It is particularly noteworthy that under OoD conditions, such as in the Advection $(\beta=1)$ and Reaction-Diffusion scenarios, the fine-tuned PDEformer rapidly attains superior results. This highlights the model’s few-shot learning [^6] In this work, we use the term few-shot learning to describe the model’s proficiency in adapting to and learning from new data that falls outside the distribution of the training set, using only a small number of examples for fine-tuning. ability in adapting to unfamiliar scenarios.
@@
Table 3: Test relative $L^{2}$ error on PDEBench, in which the PDE coefficients lie outside the range of the pretraining data. We format the first and second best outcomes in bold and underline, respectively.
Model Burgers Advection Reaction-Diffusion
$\nu=0.001$ $\beta=1$ $\nu=1,~{}\rho=1$
U-Net (Ronneberger et al., 2015) 0.2431 0.2655 0.0126
Autoregressive U-Net 0.2865 0.3735 0.0055
DeepONet (Lu et al., 2021) 0.2010 0.0187 0.0015
FNO (Li et al., 2021) 0.0700 0.0097 0.0018
PDEformer-FS (Ours) 0.0645 0.0239 0.0013
PDEformer (Ours) 0.0921 0.4000 0.7399
PDEformer-FT (Ours) 0.0295 0.0075 0.0009
@@ Appendix E Inference Time
Appendix F Autoregressive U-Net
The U-Net model exhibits unsatisfactory performance in our experiments, and some may speculate that the practice of predicting the entire spatial-temporal solution is not suitable for U-Nets. To address these concerns, we also implement an autoregressive variant of the U-Net model. Following PDEBench Takamoto et al. (2022), this model takes $\ell$ consecutive timesteps as the input, and predicts the next unknown timestep. In other words, the model approximates the mapping $[u(t-\ell,\cdot),\dots,u(t-1,\cdot)]\mapsto\hat{u}(t,\cdot)$. The model architecture is analogous to the non-autoregressive U-Net, except that it now operates on one-dimensional data.
During training, we randomly select $\ell$ consecutive timesteps from a data sample, feed it into the U-Net model, and rollout to predict the next $K$ timesteps:
$\begin{split}[u(t-\ell,\cdot),\dots,u(t-2,\cdot),u(t-1,\cdot)]&\mapsto\hat{u}(t,\cdot),\\ [u(t-\ell+1,\cdot),\dots,u(t-1,\cdot),\hat{u}(t,\cdot)]&\mapsto\hat{u}(t+1,\cdot),\\ \cdots\\ [\hat{u}(t-\ell+K-1,\cdot),\dots,\hat{u}(t+K-2,\cdot)]&\mapsto\hat{u}(t+K-1,\cdot).\end{split}$
The loss function is a weighted average of the prediction error, in the form
$\mathcal{L}(\theta)=\sum_{k=0}^{K-1}\lambda_{k}\cdot\mathrm{nRMSE}(u(t+k,\cdot),\hat{u}(t+k,\cdot)).$
In the implementation, we select $\ell=4$, $K=16$, $\lambda_{0}=1$, $\lambda_{1}=\cdots=\lambda_{15}=0.1$.
For the inference phase, we feed the first $\ell$ timesteps into the model, and rollout until we obtain the entire spatial-temporal solution. Note that all the other models involved in our experiments only takes the initial value (i.e. the first timestep) as the network input. Figure 10 illustrates the predictions of the autoregressive U-Net model. We notice that the model successfully captures the overall dynamics inside the spatial interval, and exhibits a high per-step prediction accuracy. However, small error would appear near the boundary points, and is then amplified during the rollout prediction process, leading to unsatisfactory spatial-temporal prediction results. Indeed, such boundary errors might be mitigated if we modify the network architecture to enforce periodicity, but the resulting network design would then be equation-specific, and is not applicable to more general PDEs with non-periodic boundary conditions.
Figure 10: Prediction results of the autoregressive U-Net model. Top: Burgers’ equation with $\nu=0.1$. Bottom: Advection equation with $\beta=0.1$. The horizontal axis corresponds to the spatial coordinate $x$, and the vertical axis corresponds to the temporal axis $t$.